a ball thrown up is caught after 6s.the height to which the ball has risen is

Grade:11

3 Answers

Bibhu prashad nayak
5 years ago
Hello friend..!!!
in the above question , we should calculate the height attained by the ball ,
we know H = ut - 1/2gt²
u = ?? 
t = 6 sec 
h = ??
we know that ,
v = u + at  ( but here v = -u before reaching the ground )
-u = u + at 
-2u = at  ( a = g = 9.8 and t = 6 )
u = ( 9.8 x 6 ) / 2 
u = 29 . 4 m / s -----------------------( 1 ) 
we know,
H = ut - 1/2 gt²
H = (29.4)(3) - 1/2(9.8)(3x 3 )
here i considered t = 3 sec because , the question asked is for max height so if the total time taken for the journey is 6 sec then half of it is 3 sec.
H = 44.11 m

thats it.......
AKSHANSH PAREEK
5 years ago
accleration of body  will be g=10m/sec2
total time of flight=6sec
height to which ball has risen = total distance/ 2
final velocity=0 as body comes to rest
now v=u+gt
0=-u+10*6
u=60m/sec
now v2=-u2+2gs
s=90m
height to which  ball has risen=s/2
=45m
 
 
 
AKSHANSH PAREEK
5 years ago
answere which i had given is also correct as i have taken g=10m/sec2 not 9.8m/sec2  please approve my answer that was correct